40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8547     Entry time: Tue May 7 23:03:12 2013
Author: Koji 
Type: Configuration 
Category: CDS 
Subject: CDS work 

Summary:

c1rfm / c1lsc / c1ass / c1sus were modified. They were recomplied and installed. They are running fine
and confirmed PRMI locking (attempt), arm locking, and Yarm ass with the new codes.

Motivation:

1a. SQRTing switching for POP110 was wrong. 0 enabled sqrting, 1 disabled sqrting. I wanted to fix this.
1b. Sqrting for POP22 was not implemented.

2. Preparation for the shadow sensor control with POPDC.

3. ASS had only an input. I want to run two ASS for the X and Y arms.

SQRTing for POP110/22:

- Flipped the input of the bypass switch. Correspoding MEDM indicators are fixed on the power normalization screen.
- Copied the sqrting structure from POP110 to POP22. Correspoding MEDM buttom was made on the power normalization screen.

- The function of the sqrting buttons were confirmed.

Additional ASS output:

- The output path "NPRO" was removed. Corresponding RFM channels have also removed.
- The previous NPRO path was turned to the "ASS1" path. The previous "ASS" path was turned to "ASS2".
- Corresponding shared memory channel are created/renamed.
- c1ass was modified to receive the new ASS shared memory channels. ASS1 is assigned to the X arm. ASS2 is assigned to the Y arm
- The output matrix screen and the lockin screen were modified accordingly.
- Only script/ASS/Arm_ASS_Setup.py was affected. The corespoding lines (matrix assignment) was fixed.

- The function of Den's version of  ASS was confirmed.

LSC->PRM ASC path

- We want to connect POPDC to PRM ASC. POPDC is acquired on c1lsc.
- So, for now we use the LSC input matrix to assign POPDC to one of the servo bank.
- The last row of the LSC output matrix was assigned to the PCIE connection to c1sus.
- This PCIE connection was connected to the PRM ASC YAW input.

- The connection between LSC and SUS was confirmed.

- During this process I found that there are bunch of channels transferred from LSC to SUS via RFM.
  These channels are transferred via PCIE(dolphin) and then via RFM. But LSC and SUS are connected
  with dolphin. So this just adds additional sampling delay while there is no benefit. I think we should remove the RFM part.
  Note that we need to use RFM for the end mirrors but this also should use only the RFM connection.


Rebuilding the codes

- Prior to the tests of the new functionalities, the codes were rebuild/installed as usual.
- The suspension were shutdown with the watch dogs before the restart of the realtime codes.
- Once the realtime codes were restarted successfully, the watch dogs were reloaded.
- As we removed/added the channels, fb was restarted.
- c1rfm / c1lsc / c1ass / c1sus codes were checked-in to svn
 

ELOG V3.1.3-